A+C=2B , then cosC-cosA / sinA-sinC should be equal to ?

Grade:11

1 Answers

Rajeswari
6 years ago
A+c=2b ,cosc-cosa=-2sin(c+a)/2sin(a-c)/2÷2cos(c+a)/2sin(a-c)/2. =-2sin2b/2÷2cos2b/2. =sinb/cosb. Tanb